OpenStudy (anonymous):

Factor completely. If the polynomial is not factorable, write prime. A. b^2 -16b + 16 B. 10c^2 -17c +3 C. p^2 + 2p - 24 D. c^3 +8

OpenStudy (anonymous):

@mathmath333 @amistre64 @aum @Conqueror @ilovebmth1234 @ILovePie

OpenStudy (anonymous):

@iGreen

OpenStudy (mathmath333):

what do u think ,which are factorable and which are not

OpenStudy (mathmath333):

\( c^3+8\\ \text{ is clearly factorable due to formula}\\ a^3-b^3=(a-b)(a^2+ab+b^2)\)

OpenStudy (mathmath333):

\(\text{in quafratic equation u can findout }\\ \text{if the equation is factorable or not with determinant}\\ ax^2+bx+c=0\\then\\D=b^2-4ac\\\text{if D is perfect square number then it is factorable}\)